OpenStudy (anonymous):

Which of the following are measurements of the sides of a right triangle? A. 10, 8, 6 B. 13, 12, 5 C. 26, 24, 10 D. all of the above

Parth (parthkohli):

All of them lol..

OpenStudy (anonymous):

how so?

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow a^2 + b^2 = c^2 }\) c is the longest side.

OpenStudy (anonymous):

all

Parth (parthkohli):

The formula I stated is called Pythagorean theorem.

Parth (parthkohli):

@CarmelGirlKatie47 Good work :) I like you asking 'Why'

OpenStudy (anonymous):

phrthagren therum i spelt that rong but close

Parth (parthkohli):

haha lol miranda

OpenStudy (anonymous):

A2+b2 =c2

Parth (parthkohli):

You get that carmel?

OpenStudy (anonymous):

sorry my computer was lagging. yeah i did. i just didn't understand where all the other numbers came from lol. but thank you i got it now.

OpenStudy (anonymous):

parth your a lifesaver. fanning you.

Parth (parthkohli):

Let me try it with 6,8,10 for you. \(\Large \color{MidnightBlue}{\Rightarrow 6^2 + 8^2 = 10^2 }\) \(\Large \color{MidnightBlue}{\Rightarrow 36 + 64 = 100 }\) \(\Large \color{MidnightBlue}{\Rightarrow 100 = 100 \checkmark }\)

OpenStudy (anonymous):

do they all have to equal out 100? what do they all have that are similar?

Parth (parthkohli):

No, I mean that when you set up an equation like that, the values you get are equal. If they are not, then they can't be a right triangle.

OpenStudy (anonymous):

oh i see. okay so they all have to be congruent to each other?

Parth (parthkohli):

Let's try it with 5,12,13: \(\Large \color{MidnightBlue}{\Rightarrow 5^2 + 12^2 = 13^2 }\) \(\Large \color{MidnightBlue}{\Rightarrow 25 + 144 = 169 }\) \(\Large \color{MidnightBlue}{\Rightarrow 169 = 169 }\) This tells that they can be sides of a right angled triangle.

Parth (parthkohli):

Okay, give me three numbers, I'll test if they can be sides of a right triangle.

OpenStudy (anonymous):

9, 21, 19

OpenStudy (anonymous):

are all the numbers squared?

Parth (parthkohli):

Okay, the normal equation for testing this is: \(\Large \color{MidnightBlue}{\Rightarrow a^2 + b^2 = c^2 }\) c is the largest number. If the equation is satisfied, then they can be. \(\Large \color{MidnightBlue}{\Rightarrow 9^2 + 19^2 = 21^2 }\) \(\Large \color{MidnightBlue}{\Rightarrow 81 + 361 = 441 }\) \(\Large \color{MidnightBlue}{\Rightarrow 442 \ne 441 }\) They can't be the three sides.

OpenStudy (anonymous):

do they have to be in the bounderies of 360? 30+60+90

Parth (parthkohli):

I'm not sure what you just meant by that.

Parth (parthkohli):

Just keep it simple. When you are given three numbers and asked if they can be the three sides of a right-angled triangle, then just test it with that equation.

Parth (parthkohli):

Remember that c is the largest number.

OpenStudy (anonymous):

im starting to understand this, let me try one on my own.

Parth (parthkohli):

Message me if you have any confusion :D

OpenStudy (anonymous):

okay so for this one i got none of the above. because they didn't equal with C Which of the following are measurements of the sides of a right triangle? A. 101, 99, 20 B. 28, 26, 12 C. 17, 14, 6 D. none of the above

Parth (parthkohli):

A. works